Ncert Solutions Maths class 12th

Get insights from 2.5k questions on Ncert Solutions Maths class 12th, answered by students, alumni, and experts. You may also ask and answer any question you like about Ncert Solutions Maths class 12th

  • 2.5k

    Questions

  • 0

    Discussions

  • 1

    Active Users

  • 65

    Followers

Have a question about Ncert Solutions Maths class 12th? Ask our experts

Ask Your Question

Keep it short & simple. Type complete word. Avoid abusive language.

New answer posted•a year ago
S
a year ago

83. The probability of getting a six in a throw of die is 1/6 and not getting a six is 5/6.

Let p = 1/6 and q = 5/6

The probability that the 2 sixes come in the first five throws of the die is

5C2
(16)2 (56)3=10× (5)3 (6)5

∴ Probability that third six comes in the sixth throw =

=10× (5)3 (6)5×16=10×125 (6)6=10×12546656=62523328

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

82. Let p and q respectively be the probabilities that the player will clear and knock down the hurdle.

∴P=56⇒q=1−p=1−56=16

Let X be the random variable that represents the number of times the player will knock down the hurdle.

Therefore, by binomial distribution, we obtain

P(X=x)=nCx
pn−xqx

P (player knocking down

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

81. Total number of balls in the urn = 25

Balls bearing mark ‘X’ = 10

Balls bearing mark ‘Y’ = 15

p = P (ball bearing mark ‘X’) =10/25 = 2/5

q = P (ball bearing mark ‘Y’) =15/25 = 3/5

Six balls are drawn with replacement. Therefore, the number of trials are Bernoulli trials.

Let Z be the random

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

80. It is given that 90% of people are right handed.

p=P (right handed) =90%=90100=910 and q=P (left handed) =1−910=110

Using Binomial distribution,

Probability more than 6 people are right handed is given by

=∑r=71010cr

prqn−r

=∑r=71010cr

(910)r(110)10−r

Hence, probability of having at most 6 right handed people:

=1−P (more than

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

79. Given, Men having grey hair =5%

Women having grey hair =0.25%

Total people with grey hair = (5+0.25)% =5.25%

Probability that the selected person’s hair is of male =55.25=11.05=100105=2021

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

78. i. Here, Sample space of given condition,

S={(B,B),(B,G),(G,B),(G,G)}

Let, E: denotes the event both children are male

F: denotes the event at least one of the children is a male

∴E∩F={(B,B)}

⇒P(E∩F)=14

Here, P(E)=14 and P(F)=34

⇒P(E/F)=P(E∩F)P(F)=1434=13

ii. Let, A: event that both children are female

B: event that elder child is female

A={G,G} P(A)=14

B={(G,G),(G,B)} P(B)=24=12

A∩B={(G,G)} P(A∩B)=14

P(A/B)=P(A∩B)P(B)

=1412=14×2=12

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

77. A is a subset of B

⇒A∩B=A

∴P (A∩B)=P (B∩A)=P (A)

=P (B/A)=P (B∩A)P (A)=1

A∩B≠0

(A∩B)=0

⇒P (B/A)=P (B∩A)P (A)=0

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

76. The repeated tossing of a die are Bernoulli trials. Let X represent the number of times of getting 5 in 7 throws of the die.

Probability of getting 5 in a single throw of the die, p = 1/6

∴q=1−p=1−16=56

Clearly, X has the probability distribution with n=7 and P=16

∴P(X=x)=nCx
qn−xpx=7Cx
(56)7−x.(16)x

P (getting 5 exactly twice) =P(X=2)

=7C2
(56)5.(16)2=21.(56)5.136=(712)+(56)5

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

75. Let X represent the number of winning prizes in 50 lotteries. The trials are Bernoulli trials.

Clearly, X has a binomial distribution with n = 50 and p = 1/100

∴q=1−p=1−1100=99100∴P(X=x)=nCx
qn−xpx=50Cx
(99100)50−x.(1100)x

P (winning at least once) =P(X≥1)

=1−P(X<1)=1−P(X=1)=1−50C0

(99100)50=1−1.(99100)50=1−(99100)50

P (winning exactly once) =P(X=1)

=50C1

(99100)49.(1100)1=50(1100)(99100)49=12(99100)49

P (at least twice) =P(X≥2)

=1−P(X<2)=1−P(X≤1)=1−[P(X=0)+P(X=1)]=[1−P(X=0)]−P(X=1)=1−(99100)50−12.(99100)49=1−(99100)49.[99100+12]=1−(99100)49.(149100)=1−(149100)(99100)49

...Read more

Thumbs Up IconUpvote Thumbs Down Icon
New answer posted•a year ago
S
a year ago

74. The repeated guessing of correct answers from multiple choice questions are Bernoulli trials. Let X represent the number of correct answers by guessing in the set of 5 multiple choice questions.

Probability of getting a correct answer is, p = 1/3

∴q=1−p=1−13=23

Clearly, X has a binomial distribution wi

...Read more

Thumbs Up IconUpvote Thumbs Down Icon